# Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity

### v1

@@ A.2.3 Self-Enhancement
Recent LAIF research utilizes outputs of more capable models compared to the policy model. For instance, Sharma et al. (2024a) utilizes dialogue outputs from GPT-3.5 and GPT-4 to fine-tune smaller SFT models. Also, the chosen and rejected outputs are dependent on the larger models. Lee et al. (2024) attempted to train using policies and AI feedback models of the same size (e.g., PaLM 2 XS (Anil et al., 2023)). However, as the exact model sizes are not disclosed, it is challenging to estimate the scalability of same-sized LAIF setups.
We train LAIF on our binary QA task (i.e., GPQA and QuALITY). In typical setups, the same dialogue and model outputs are ranked in order, and chosen and rejected outputs are created for every $\binom{K}{2}$ combination, constructing a feedback dataset. Human preference data is usually composed of around four output results (Ouyang et al., 2022b).
@@
A.2.4 Weak-to-Strong Generalization
@@
For the noise test, we examine the performance degradation caused by noise generated by weakly fine-tuned models across varying AI system sizes. Our study utilizes GPT-2 variants (Base, Medium, and Large), Phi-3.5 Mini-Instruct, and Qwen2.5 models (3B and 7B Instruct). By testing the impact of noise, we aim to reveal overfitting issues in large-parameter models and their effects on performance. We use the full SciQ (Welbl et al., 2017) and CosmosQA (Huang et al., 2019) training and test sets. We format the input as shown in Tab. A.4. To induce noise, we randomly flip a certain percentage of the original labels in the training set, varying from 0% (0.1) to 100% (1.0). We then measure both training accuracy and test accuracy. For W2SG training, we follow the setup from the original work (Burns et al., 2023).
For multi-objective alignment, the current W2SG approach is mainly limited to binary classification tasks, such as question answering or reward modeling. Recent work (Yang et al., 2024b) investigated training W2SG models using DPO (Rafailov et al., 2023) and SimPO (Meng et al., 2024) alignment methods, comparing their performance. However, the findings showed that in realistic multi-objective alignment scenarios, such as Anthropic-HHH (Bai et al., 2022a) or Ultrafeedback (Cui et al., 2024), W2SG fails to fully achieve its intended goal—where a strong, fine-tuned student model consistently outperforms a weak, fine-tuned teacher model when evaluated with reward accuracy. Recently, the RewardBench benchmark (Lambert et al., 2024) is introduced, enabling preference evaluation across challenging, diverse, and out-of-distribution query domains, including chat, reasoning, and safety. We utilize RewardBench to investigate the feasibility of multi-objective alignment in W2SG, at least in a preliminary form. For this, we use Qwen2.5-0.5B-Instruct as the weak AI system and Qwen-1.5B-Instruct as the strong AI system.
@@
A.3 Experiment Results
A.3.1 Sandwiching
@@
A.3.2 Self-Enhancement
Self-Enhancement aims to evaluate self-improvement ability using LAIF with DPO training, especially when supervision comes from equally capable models. The result is shown in Tab. A.4. The table shows performance (i.e., accuracy) with GPQA and QuALITY benchmarks on three different model sizes on Qwen2.5-Instruct 0.5B, 3B, and 7B.
@@ A.3.2 Self-Enhancement
‘LAIF--SFT’ represents the performance difference between the LAIF model and the SFT model. The results clearly demonstrate that training with self-generated data can negatively impact performance. For instance, for the 0.5B model, LAIF--SFT shows an improvement of 2.15 and 0.9 in GPQA and QuALITY, respectively. However, for the 7B model, LAIF--SFT shows a decrease in performance by -2.67 and -7.99.
A.3.3 Weak-to-Strong Generalization
@@
In multi-objective alignment, we examine the feasibility of multi-objective preference alignment using UltraFeedback dataset (Cui et al., 2024) which aims to align a model to multiple objectives simultaneously. We use RewardBench (Lambert et al., 2024) to evaluate performance across four aspects: chat, chat hard, safety, and reasoning. Fig. LABEL:fig:rewardbench_evaluation_all presents the aggregated RewardBench scores in the W2SG setup, using Qwen2.5-0.5B and Qwen2.5-1.5B instruction-tuned models as base models.
@@
A.4 Debate Argument Analysis
@@
In the first round of the debate, the main difference is in their approaches to evidence. Debater 1 focuses on explicit textual evidence, citing specific quotes from the story. In contrast, Debater 2 relies heavily on inference. Debater 1 argues that the story describes only the return trip from Callisto to Earth, citing the lack of evidence for a round trip. Debater 2, however, emphasizes the word “back,” inferring that it implies a round trip.
In the second round, Debater 1 defends its position by challenging Debater 2’s reliance on implicit assumptions and the absence of explicit evidence. Debater 2 counters by asserting that the term “back” inherently implies a round trip but fails to provide additional textual evidence to substantiate this claim.
In the third round, Debater 1 highlights that Debater 2’s argument relies on assumptions rather than direct evidence. Debater 2 continues to argue that the implied meaning of “back” is sufficient to infer a round trip but does not address the lack of explicit confirmation in the story.
In short, Debater 1’s argument is more convincing to the judge because it is grounded in explicit textual evidence and direct quotes from the story. In contrast, Debater 2’s argument relies on assumptions about the term “back” without supporting textual evidence.
@@
Table 11: Datasets used for LAIF evaluation. Both datasets are considered challenging enough to evaluate scalable oversight. For GPQA reasoning, we use the main set recommended by the authors for primary experiments. For QuALITY long reading comprehension, we filter the dataset using the same strategy as in (Khan et al., 2024) and divide it into two subsets: one for training and development, and the other for testing. The numbers in parentheses represent the number of unique stories in each subset, which are not shared across subsets.

### v2

@@ A.7.3 Self-Enhancement
Recent LAIF research utilizes outputs of more capable models compared to the policy model. For instance, Sharma et al. [161] utilizes dialogue outputs from GPT-3.5 and GPT-4 to fine-tune smaller SFT models. Also, the chosen and rejected outputs are dependent on the larger models. Lee et al. [105] attempted to train using policies and AI feedback models of the same size (e.g., PaLM 2 XS [8]). However, as the exact model sizes are not disclosed, it is challenging to estimate the scalability of same-sized LAIF setups.
We train LAIF on our binary QA task (i.e., GPQA and QuALITY). In typical setups, the same dialogue and model outputs are ranked in order, and chosen and rejected outputs are created for every $\binom{K}{2}$ combination, constructing a feedback dataset. Human preference data is usually composed of around four output results [135].
@@
A.7.4 Weak-to-Strong Generalization
@@
For the noise test, we examine the performance degradation caused by noise generated by weakly fine-tuned models across varying AI system sizes. Our study utilizes GPT-2 variants (Base, Medium, and Large), Phi-3.5 Mini-Instruct, and Qwen2.5 models (3B and 7B Instruct). By testing the impact of noise, we aim to reveal overfitting issues in large-parameter models and their effects on performance. We use the full SciQ [189] and CosmosQA [84] training and test sets. We format the input as shown in Tab. A.9. To induce noise, we randomly flip a certain percentage of the original labels in the training set, varying from 0% (0.1) to 100% (1.0). We then measure both training accuracy and test accuracy. For W2SG training, we follow the setup from the original work [31].
For multi-objective alignment, the current W2SG approach is mainly limited to binary classification tasks, such as question answering or reward modeling. Recent work [198] investigated training W2SG models using DPO [148] and SimPO [119] alignment methods, comparing their performance. However, the findings showed that in realistic multi-objective alignment scenarios, such as Anthropic-HHH [14] or Ultrafeedback [46], W2SG fails to fully achieve its intended goal—where a strong, fine-tuned student model consistently outperforms a weak, fine-tuned teacher model when evaluated with reward accuracy. Recently, the RewardBench benchmark [103] is introduced, enabling preference evaluation across challenging, diverse, and out-of-distribution query domains, including chat, reasoning, and safety. We utilize RewardBench to investigate the feasibility of multi-objective alignment in W2SG, at least in a preliminary form. For this, we use Qwen2.5-0.5B-Instruct as the weak AI system and Qwen-1.5B-Instruct as the strong AI system.
@@
A.8 Experiment Results
A.8.1 Sandwiching
@@
A.8.2 Self-Enhancement
Self-Enhancement aims to evaluate self-improvement ability using LAIF with DPO training, especially when supervision comes from equally capable models. The result is shown in Tab. LABEL:tab:aif_result. The table shows performance (i.e., accuracy) with GPQA and QuALITY benchmarks on three different model sizes on Qwen2.5-Instruct 0.5B, 3B, and 7B.
@@ A.8.2 Self-Enhancement
‘LAIF-SFT’ represents the performance difference between the LAIF model and the SFT model. The results clearly demonstrate that training with self-generated data can negatively impact performance. For instance, for the 0.5B model, LAIF-SFT shows an improvement of 2.15 and 0.9 in GPQA and QuALITY, respectively. However, for the 7B model, LAIF-SFT shows a decrease in performance by -2.67 and -7.99.
A.8.3 Weak-to-Strong Generalization
@@
In multi-objective alignment, we examine the feasibility of multi-objective preference alignment using UltraFeedback dataset [46] which aims to align a model to multiple objectives simultaneously. We use RewardBench [103] to evaluate performance across four aspects: chat, chat hard, safety, and reasoning. Fig. LABEL:fig:rewardbench_evaluation_all presents the aggregated RewardBench scores in the W2SG setup, using Qwen2.5-0.5B and Qwen2.5-1.5B instruction-tuned models as base models.
@@
A.9 Debate Argument Analysis
@@
In the first round of the debate, the main difference is in their approaches to evidence. Debater 1 focuses on explicit textual evidence, citing specific quotes from the story. In contrast, Debater 2 relies heavily on inference. Debater 1 argues that the story describes only the return trip from Callisto to Earth, citing the lack of evidence for a round trip. Debater 2, however, emphasizes the word “back," inferring that it implies a round trip.
In the second round, Debater 1 defends its position by challenging Debater 2’s reliance on implicit assumptions and the absence of explicit evidence. Debater 2 counters by asserting that the term “back" inherently implies a round trip but fails to provide additional textual evidence to substantiate this claim.
In the third round, Debater 1 highlights that Debater 2’s argument relies on assumptions rather than direct evidence. Debater 2 continues to argue that the implied meaning of “back" is sufficient to infer a round trip but does not address the lack of explicit confirmation in the story.
In short, Debater 1’s argument is more convincing to the judge because it is grounded in explicit textual evidence and direct quotes from the story. In contrast, Debater 2’s argument relies on assumptions about the term “back" without supporting textual evidence.
@@
Debate Judge Prompt
You are an expert judge in a structured debate where two participants are arguing over the correct reasoning and answer to a math-related question. Your task is to determine which participant provides the most convincing argument based on logic, mathematical correctness, and clarity of explanation. Note that the correctness of the arguments must align with the mathematical reasoning provided in the debate.
Here is the question and the two possible answers:
Question: [Question]
A: [Answer A]
B: [Answer B]
Participant A is advocating for answer A, and Participant B is advocating for answer B.
Here’s the current transcript of the debate:
<transcript>[Transcript]</transcript>
When evaluating the debate, focus on these criteria: 1. Logical Consistency: Are the arguments logically sound and free of contradictions? 2. Mathematical Rigor: Do the participants demonstrate correct and precise mathematical reasoning? 3. Clarity: Are the arguments clearly explained and easy to follow?
Think about your evaluation in <thinking></thinking> tags by following these steps: 1. Analyze the strengths and weaknesses of each participant’s arguments. 2. Identify any logical or mathematical errors. 3. Decide step by step which argument is most compelling and aligns with the correct mathematical reasoning. On a new line, format your final answer as “Answer: <A|B>”.
Table 11: Datasets used for LAIF evaluation. Both datasets are considered challenging enough to evaluate scalable oversight. For GPQA reasoning, we use the main set recommended by the authors for primary experiments. For QuALITY long reading comprehension, we filter the dataset using the same strategy as in [95] and divide it into two subsets: one for training and development, and the other for testing. The numbers in parentheses represent the number of unique stories in each subset, which are not shared across subsets.
@@
Figure 7: Logical structure of our position. We argue that (a) superalignment is achievable, and (b) research on it must begin now, supported by a formal definition. Our proposed conceptual framework (c) is grounded in the joint optimization of task competence and value conformity, based on two principles: determining the capability–capacity gap and increasing signal diversity.